OpenStudy (anonymous):

let f(x)=cubed root of x If a does not =0 find f'(a) using the definition of a derivative.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

definition, not the power rule right?

OpenStudy (anonymous):

yeah, I am having some trouble because of the cubed root

OpenStudy (misty1212):

i bet i can show you the gimmick did you do it with the square root ever?

OpenStudy (anonymous):

yeah I have

OpenStudy (misty1212):

ok the idea with the square root is to multiply by the conjugate, because \[a^2-b^2=(a+b)(a-b)\] so the radical goes away

OpenStudy (misty1212):

but for the cubed root you have to use the difference of two cubes, not the difference of two squares that will get rid of the cubed root i.e. \[a^3-b^3=(a-b)( a^2+ab+b^2)\]

OpenStudy (anonymous):

ohhh I was doing it as though it was a squared .-.

OpenStudy (anonymous):

okay so how exactly will that look for this problem

OpenStudy (misty1212):

\[\frac{\sqrt[3]{a+h}-\sqrt[3]{a}}{h}\] multiply top and bottom by \[\sqrt[3]{(a+h)^2}+\sqrt[3]{a}\sqrt[3]{a+h}+\sqrt[3]{a^2}\]

OpenStudy (misty1212):

don't really multiply it out the numerator will be \[a+h-a=h\]

OpenStudy (misty1212):

leave the denominator in factored form cancel the \(h\) top and bottom and then replace \(h\) by \(0\)

OpenStudy (misty1212):

you are using this \[(a-b)(a^2+ab+b^2)=a^3-b^3\]with \[a=\sqrt[3]{a+h},b=\sqrt[3]{a}\]

OpenStudy (anonymous):

I am still a a but confused..

OpenStudy (anonymous):

*a bit

OpenStudy (anonymous):

@misty1212